### v3

# Detection of evolutionary shifts in variance under an Ornstein–Uhlenbeck model

Abstract
Sudden changes in environmental conditions can lead to evolutionary shifts not only in the optimal trait value ($\theta$), but also in the diffusion variance ($\sigma^{2}$) under the Ornstein-Uhlenbeck (OU) model. While several methods have been developed to detect shifts in optimal values, few explicitly account for concurrent shifts in both evolutionary variance and diffusion variance. We use a multi-optima and multi-variance OU model to describe trait evolution with shifts in both $\theta$ and $\sigma^{2}$ and analyze how covariance between species is affected when shifts in variance occur along the phylogeny. We propose a new method that simultaneously detects shifts in both variance and optimal values by formulating the problem as a variable selection task using an $\ell_{1}$-penalized loss function. Our method is implemented in the R package ShiVa (Detection of evolutionary shifts in variance). Through simulations, we compare ShiVa with existing methods that can automatically detect evolutionary shifts under the OU model ($\ell$1ou, PhylogeneticEM, and PCMFit). Our method demonstrates improved predictive ability and significantly reduces false positives in detecting optimal value shifts when variance shifts are present. When only shifts in optimal value occur, our method performs comparably to existing approaches. We apply ShiVa to empirical data on floral diameter in Euphorbiaceae and buccal morphology in Centrarchidae sunfishes.
Key words: evolutionary shift detection, Ornstein-Uhlenbeck model, LASSO, trait evolution, phylogenetic comparative methods
1 Introduction
The process of trait evolution is influenced by environmental conditions. When species experience abrupt environmental changes, their evolutionary processes may also change to adapt to the new environment (Losos,, 2011; Mahler et al., 2013b, ). These changes in the evolutionary process leave signatures in the observed traits of present-day species. By detecting such evolutionary shifts based on observed trait data, we gain valuable insights into historical environmental changes and the evolutionary history of species.
We typically model evolutionary shifts as changes in the value of the parameters of trait evolution models which assume that traits evolve along a phylogenetic tree according to a continuous time Markov process. The two most popular trait evolution models are the Brownian Motion (BM) model (Felsenstein,, 1985) and the Ornstein-Uhlenbeck (OU) model (Hansen,, 1997). This paper will focus on detecting evolutionary shifts under the OU model, which considers both neural drift and natural selection. Although several researchers have studied this topic, much of the existing work focuses only on detecting shifts in the optimal value of the model (Butler and King,, 2004; Uyeda and Harmon,, 2014; Ho and Ané, 2014a, ; Khabbazian et al.,, 2016; Bastide et al.,, 2017; Zhang et al.,, 2024). In reality, abrupt environmental changes are likely to influence not only the optimal value but also the diffusion variance, reflecting how much random fluctuation occurs around the evolutionary process.
In this paper, the term variance specifically refers to the diffusion variance ($\sigma^{2}$) of the OU model, which quantifies the background rate of stochastic variation in trait evolution. This variance may increase in unstable environments or decrease under stabilizing conditions. Some existing methods allow modeling shifts in evolutionary parameters, including the optimal value ($\theta$), the rate of adaptation ($\alpha$), and the evolutionary variance ($\sigma^{2}$). Pagel and Meade, (2006) introduced BayesTraits, a Bayesian framework that models evolutionary shifts across predefined regimes, supporting changes in both optimal trait values and evolutionary rates (modeled through $\alpha$). Beaulieu et al., (2012) developed the OUwie package, which fits a suite of OU-based models that allow for changes in $\theta$, $\alpha$, and $\sigma^{2}$ between user-specified regimes, using maximum likelihood estimation. Clavel et al., (2015) extended these ideas to multivariate traits with mvMorph, fitting multivariate Brownian motion and OU models that accommodate shifts in evolutionary rates and covariances across lineages, also requiring predefined regime mappings. More recently, Gaboriau et al., (2020) developed JIVE, a multi-platform R package for analyzing both intra- and interspecific trait evolution, allowing flexible modeling of shifts in evolutionary parameters, including diffusion variance, under both continuous and discrete regimes. Bastide et al., (2021) proposed a Bayesian framework that models shifts in evolutionary variance without requiring prior regime specification, using Hamiltonian Monte Carlo (HMC) sampling to jointly infer branch-specific parameters and variance shifts. Several methods specifically aim to detect shifts in diffusion variance under the BM model. Eastman et al., (2011) proposed a Bayesian relaxed BM framework, where branch-specific evolutionary rates ($\sigma^{2}$) are drawn from a lognormal distribution. Grundler et al., (2022) introduced a maximum-likelihood dynamic programming approach to efficiently infer branch-specific rates. For the OU model, Mitov et al., (2019) introduced the PCMFit package, which, to our knowledge, is the only available method capable of automatically detecting shifts in diffusion variance ($\sigma^{2}$) without predefined regimes. However, PCMFit assumes that all model parameters shift together across the same set of regimes. In contrast, our method allows independent shifts in the optimal value ($\theta$) and the diffusion variance ($\sigma^{2}$), providing modeling flexibility by permitting shifts to occur at different locations on the tree.
While these tools allow evolutionary shifts to be modeled, they typically require either pre-specified regimes or rely on computationally intensive full Bayesian inference. Furthermore, few methods explicitly focus on detecting both shifts in optimal value and shifts in variance simultaneously using a data-driven variable selection framework. In this paper, we propose a novel LASSO-based method that detects both types of shifts within a unified variable selection framework. Our method uses a multi-optima and multi-variance OU process model, similar to the models used in OUwie (Beaulieu et al.,, 2012), but innovates by formulating the detection of both shifts in optima and shifts in variance as a single variable selection problem. This approach allows flexible and efficient detection of shifts across the entire tree, without requiring a priori specification of regime boundaries. We implement this approach in a new R package, ShiVa (Detection of Evolutionary Shifts in Variance), which detects both shifts in optimal values and variance under the Ornstein–Uhlenbeck model for ultrametric trees and univariate traits. Through simulation studies, we compare ShiVa to existing methods, including $\ell$1ou, PhylogeneticEM and PCMFit evaluating performance in terms of both shift detection accuracy and predictive performance.
The remainder of the paper is organized as follows. The Trait evolution with shifts in both optimal value and variance section introduces the trait evolution model incorporating shifts in both optimal value and variance, and formally defines the shift detection problem. The Methods section describes the LASSO-based shift detection algorithm. The Simulations section presents simulation results, including comparisons with existing methods. The Case study section illustrates the application of ShiVa to two empirical datasets: floral diameter in the parasitic plant family Euphorbiaceae using the flowerTree phylogeny (Davis et al.,, 2007), and buccal morphology in Centrarchidae sunfishes (Revell and Collar,, 2009).
2 Trait evolution with shifts in both optimal value and variance
2.1 Trait evolution models
The phylogenetic tree is reconstructed from DNA sequences and is assumed to be known in this paper. The phylogenetic tree reveals the correlation structure between trait values of different species. The trait values are correlated based on the shared evolutionary history of species. The trait values of internal nodes are hidden and only trait values of tip nodes can be observed. We assume the phylogenetic tree is ultrametric and the selection force $\alpha$ is fixed throughout the tree. When $\alpha$ is unknown, a reasonable estimate can be obtained by fitting a null (no-shift) model using the phylolm package (Ho and Ané, 2014b, ).
Trait evolution models are used to model how the trait values change over time. Brownian Motion and Ornstein–Uhlenbeck are two commonly used models to model the evolution of continuous traits. For both models, we let $Y$ denote the vector of observed trait values at the tips, and $Y_{i}$ denote the trait value of taxon $i$. For a single branch, we let $Y_{i}(t)$ denote the trait value at time $t$. These two models assume that conditioning on the trait value of a parent, the evolutionary processes of sister species are independent. We therefore only need to specify the model on one branch. For specifying the model on a single branch, we let $Y(t)$ denote the trait value at time $t$ on a fixed branch.
Brownian Motion Model
Brownian motion (BM) was first applied to model the evolution of continuous traits over time by Felsenstein, (1985). Under this model, the trait value evolves following a BM process. The process can be written as a stochastic differential equation:
$dX_{t}=\sigma(t)\,dW_{t},$
[…]
Ornstein–Uhlenbeck Model
Hansen, (1997) uses an OU process to model the evolutionary process. A selection force that pulls the trait value toward a selective optimum $\theta$ is included in OU models. An OU process $Y(t)$ is defined by the following stochastic differential equation
$dY(t)=\alpha[\theta(t)-Y(t)]dt+\sigma(t)dB(t)$
[…]
2.2 Evolutionary shifts in optimal value and variance
Butler and King, (2004) formulate the multi-optima OU model for adaptive evolution. This model assumes that the optimal value $\theta(t)$ is constant along a branch and may be different between branches. An abrupt change in $\theta(t)$ on a branch is considered an evolutionary shift in optimal value. However, in most previous work, it is assumed that the variance $\sigma^{2}(t)$ is constant throughout the tree. In reality, when a change in the environment happens, not only the optimal values, but also the variances are likely to change. In this paper, we use a multi-optima multi-variance model (Beaulieu et al.,, 2012) which allows the variances $\sigma(t)$ to change over different branches. We first derive the formulas assuming a fixed root, and later extend them to account for a random root.
Solving the OU process equation, the trait value at time $t$ is given by
[…]
Solving this, the trait value of species $i$ follows a normal distribution with expectation given by:
[…]
By Itô isometry, the element $\mathbf{\Sigma}_{i,j}$ (covariance of $Y_{i}$ and $Y_{j}$) of covariance matrix $\mathbf{\Sigma}$ is given by
[…]
Where $a$ is the separating point of species $i$ and $j$. For ultrametric trees, $t_{a}$ is the time from the root to the separating point. The distance between i and j is $2t_{ai}=2t_{aj}=2(T-t_{a})$. For shifts in optimal value, Ho and Ané, 2014a showed that the exact location and number of shifts on the same branch are unidentifiable. We therefore also assume that the optimal value is constant along a branch. We let $\theta_{b}$ denote the optimal value on branch $b$, $\Delta\theta_{b}$ denote the change in optimal value on branch $b$ and $\operatorname{path}(\textrm{root},i)$ denote the branches on the path of the phylogenetic tree from the root to the tip $i$. The expectation can be writen as the following equation (Khabbazian et al.,, 2016).
$E(Y_{i})=y_{0}e^{-\alpha T}+(1-e^{-\alpha T})\theta_{0}+\sum_{b\in\operatorname{path}(\textrm{root},i)}(1-e^{-\alpha(T-t_{b})})\Delta\theta_{b}$
Let $\beta_{0}=y_{0}e^{-\alpha T}+(1-e^{-\alpha T})\theta_{0}$ and $\beta_{b}=(1-e^{-\alpha(T-t_{b})})\Delta\theta_{b}$. Then,
$E(Y_{i})=\beta_{0}+\sum_{b}\beta_{b}\mathbf{(X_{b})}_{i}$ (1)
where $\mathbf{X_{b}}$ is a vector defined by $\mathbf{(X_{b})}_{i}=0$ if taxon $i$ is not under branch $b$, and $\mathbf{(X_{b})}_{i}=1$ if the taxon $i$ is under branch $b$, and $t_{b}$ denotes the time from the root to the beginning of branch $b$.
While shifts in the optimal value ($\theta$) affect only the expectation of each trait $Y_{i}$, shifts in variance ($\sigma^{2}$) affect both the variance of $Y_{i}$ and the covariance between traits $Y_{i}$ and $Y_{j}$. This makes the location and magnitude of variance shifts potentially more identifiable from the data than shifts in optima. Suppose species $i$ and $j$ diverged at time $t_{a}$, and along their shared ancestry there are variance shifts at times $t_{1},\ldots,t_{m}$ with magnitudes $\Delta\sigma_{1}^{2},\ldots,\Delta\sigma_{m}^{2}$. Then their covariance $\Sigma_{ij}$ is given by:
$\mathbf{\Sigma}_{ij}=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left((1-e^{-2\alpha t_{a}})\sigma_{0}^{2}+\sum_{l=1}^{m}\left(1-e^{-2\alpha(t_{a}-t_{l})}\right)\Delta\sigma_{l}^{2}\right)$
This expression shows that each variance shift contributes to the total covariance based on its position along the evolutionary path. Therefore, the timing of variance shifts affects the likelihood, and different shift placements can result in different covariance structures. This differs from the case of shifts in optima, where the shift location along a branch is not identifiable (Ho and Ané, 2014a, ). In contrast, variance shifts can, in some cases, be localized — particularly when we have multiple descendant species whose pairwise covariances respond differently depending on when the shift occurred.
To illustrate this, consider the tree below, where species $i$, $j$, $k$, and $l$ descend from a common ancestor $a$. Let there be a single variance shift on the branch from $a$ to $a^{\prime}$:
Then we can derive the following two differences in covariances:
$\mathbf{\Sigma}_{jk}-\mathbf{\Sigma}_{ij}$ $=\frac{1}{2\alpha}\left[\sigma_{0}^{2}\left(e^{-2\alpha(T-t_{a^{\prime}})}-e^{-2\alpha(T-t_{a})}\right)+\Delta\sigma^{2}\left(e^{-2\alpha(T-t_{a^{\prime}})}-e^{-2\alpha(T-t)}\right)\right]$
$\mathbf{\Sigma}_{kl}-\mathbf{\Sigma}_{ij}$ $=\frac{1}{2\alpha}\left[\sigma_{0}^{2}\left(e^{-2\alpha(T-t_{a^{\prime\prime}})}-e^{-2\alpha(T-t_{a})}\right)+\Delta\sigma^{2}\left(e^{-2\alpha(T-t_{a^{\prime\prime}})}-e^{-2\alpha(T-t)}\right)\right].$
These two equations are linearly independent and can be solved for both the shift time $t$ and the shift magnitude $\Delta\sigma^{2}$. Thus, in this configuration, the position of a single variance shift along a branch is identifiable.
However, identifiability breaks down when multiple variance shifts occur along the same branch. Suppose two variance shifts occur at times $t_{1}$ and $t_{2}$ on branch $b$, with magnitudes $\Delta\sigma_{1}^{2}$ and $\Delta\sigma_{2}^{2}$. The contribution to the covariance between any two species descending from branch $b$ will be:
$\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left[\left(1-e^{-2\alpha(t_{a}-t_{1})}\right)\Delta\sigma_{1}^{2}+\left(1-e^{-2\alpha(t_{a}-t_{2})}\right)\Delta\sigma_{2}^{2}\right].$
This combined expression depends only on the sum of contributions from both shifts. Therefore, it is not possible to uniquely identify the individual times or magnitudes of these shifts. In fact, a single shift with an appropriately chosen time and magnitude might reproduce the same covariance effect, making the two-shift model unidentifiable.
For model and computational simplicity, we assume that $\sigma^{2}$ is constant over a branch like the assumption for $\theta$. That is, we assume that any shift in variance on a branch occurs at the beginning of that branch. For shifts in mean, this assumption did not limit the space of possible models due to the unidentifiability. However, for shifts in variance, this does restrict the space of models for the trait values at the leaves, and could therefore adversely affect shift detection. We conducted simulation studies to assess the impact of these assumption violations and found that they generally do not affect detection performance, except in the case of two opposing shifts on the same branch—an edge case where the assumptions matter more. Overall, these assumptions are not a major concern. The corresponding simulation results are presented in the Supplementary Materials.
Under our assumptions, we let $\sigma^{2}_{b}$ denote the variance on the branch $b$, $\Delta\sigma^{2}_{b}$ denote the magnitude of shift in variance on the branch $b$: if no shift occurs on the branch $b$, $\Delta\sigma^{2}_{b}=0$. The covariance between species $i$ and $j$ is given by:
$\mathbf{\Sigma}_{i,j}$ $=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left((1-e^{-2\alpha t_{a}})\sigma^{2}_{0}+\sum_{b\in\operatorname{path}(\textrm{root},a)}(1-e^{-2\alpha\left(t_{a}-t_{b}\right)})\Delta\sigma^{2}_{b}\right)$
$=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left((1-e^{-2\alpha t_{a}})\sigma^{2}_{0}+\sum_{b\in\operatorname{path}(\textrm{root},a)}\left(1-e^{-2\alpha t_{a}}+e^{-2\alpha t_{a}}-e^{-2\alpha\left(t_{a}-t_{b}\right)}\right)\Delta\sigma^{2}_{b}\right)$
$=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left((1-e^{-2\alpha t_{a}})\sigma^{2}_{0}+\sum_{b\in\operatorname{path}(\textrm{root},a)}\left((1-e^{-2\alpha t_{a}})+e^{-2\alpha t_{a}}(1-e^{2\alpha t_{b}})\right)\Delta\sigma^{2}_{b}\right)$
$=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left(1-e^{-2\alpha t_{a}}\right)\sigma^{2}_{0}+\sum_{b\in\operatorname{path}(\textrm{root},a)}\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}(1-e^{-2\alpha t_{a}})\Delta\sigma^{2}_{b}$
$\qquad+\sum_{b\in\operatorname{path}(\textrm{root},a)}\frac{e^{-2\alpha T}}{2\alpha}\left(1-e^{2\alpha t_{b}}\right)\Delta\sigma^{2}_{b}.$ (2)
For the ancestral state at the root node, two different assumptions are commonly used, fixed value or stationary distribution. We assume that the ancestral state at the root is a fixed value for the above process. Equation 2.2 is the covariance of node $i$ and $j$ for the OU model with fixed root. For the OU model, the ancestral state at the root is also often assumed to have the stationary distribution. In this case, the variance of the root node is $\sigma_{0}^{2}/(2\alpha)$ (Ho and Ané,, 2013). In this case, the covariance with shifts can be written as:
$\mathbf{\Sigma}_{i,j}$ $=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left(1-e^{-2\alpha t_{a}}\right)\sigma^{2}_{0}+\sum_{b\in\operatorname{path}(\textrm{root},a)}\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}(1-e^{-2\alpha t_{a}})\Delta\sigma^{2}_{b}$
$\qquad+\sum_{b\in\operatorname{path}(\textrm{root},a)}\frac{e^{-2\alpha T}}{2\alpha}\left(1-e^{2\alpha t_{b}}\right)\Delta\sigma^{2}_{b}+\frac{e^{-2\alpha T}}{2\alpha}\sigma_{0}^{2}$
$=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\sigma^{2}_{0}+\sum_{b\in\operatorname{path}(\textrm{root},a)}\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}(1-e^{-2\alpha t_{a}})\Delta\sigma^{2}_{b}$
$\qquad+\sum_{b\in\operatorname{path}(\textrm{root},a)}\frac{e^{-2\alpha T}}{2\alpha}\left(1-e^{2\alpha t_{b}}\right)\Delta\sigma^{2}_{b}$
Let ${\gamma}_{b}=\Delta\sigma_{b}^{2}$, $\mathbf{V}_{i,j}=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}(1-e^{-2\alpha t_{a}})$, $\mathbf{U}_{i,j}=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}$, $q_{b}=\frac{e^{-2\alpha T}}{2\alpha}\left(1-e^{2\alpha t_{b}}\right)$; ${\gamma}_{b}$ is the shift in variance on branch $b$, $\mathbf{V}$ is the phylogenetic covariance matrix when $\sigma^{2}=1$ (no shift in variance) with fixed root, $\mathbf{U}$ is the phylogenetic covariance matrix when $\sigma^{2}=1$ (no shift in variance) with stationary distributed root. $\mathbf{U}$ and $\mathbf{V}$ only depend on the phylogeny and $\alpha$. The covariance between species $i$ and $j$ can be expressed as the following equation.
$\mathbf{\Sigma}_{i,j}=\begin{cases}\mathbf{V}_{i,j}\sigma_{0}^{2}+\sum_{b}{\gamma}_{b}\mathbf{X}_{ib}\mathbf{X}_{jb}\mathbf{V}_{i,j}+\sum_{b}{\gamma}_{b}q_{b}\mathbf{X}_{ib}\mathbf{X}_{jb}&\text{OU model with fixed root}\\ \mathbf{U}_{i,j}\sigma_{0}^{2}+\sum_{b}{\gamma}_{b}\mathbf{X}_{ib}\mathbf{X}_{jb}\mathbf{V}_{i,j}+\sum_{b}{\gamma}_{b}q_{b}\mathbf{X}_{ib}\mathbf{X}_{jb}&\text{OU model with random root}\end{cases}$
It consists of 3 terms. The first term is the original covariance without any shift. The last 2 terms show the influence of shifts in variance. In addition to the shift sizes $\gamma_{b}$, the second term is influenced by the phylogentic structure ($\mathbf{V}_{i,j}$), higher original covariance between two species leads to larger change in covariance for a fixed shift size (Fig. 2); the third term is influenced by the start time of branch b ($q_{b}$), earlier shifts lead to larger change in covariance between two species (Fig. 2).
Figure 1: Illustration of how original covariance affects the impact of a variance shift. The green rectangle indicates the location of a variance shift of size $\gamma_{b}$. The asterisks mark the two species whose covariance change is being evaluated. In the left panel, these two species have a higher original covariance (i.e., if no shift occurred), resulting in a larger change in covariance under the same shift, compared to the right panel.
Figure 2: Illustration of how the timing of a variance shift affects covariance change. A shift of size $\gamma_{b}$ is again marked by a green rectangle. The two asterisk-marked species share the same relationship in both panels, but in the left panel, the shift occurs earlier in their shared evolutionary history, leading to a larger change in their covariance than the same shift applied later, as in the right panel.
For simplicity, we let $\mathbf{R}$ denote the phylogenetic covariance matrix when there is no shift in variance and $\sigma^{2}_{0}=1$. So $\mathbf{R=V}$ when the root is fixed and $\mathbf{R=U}$ when the trait value at the root follows the stationary distribution.
Using matrix notation, the covariance matrix can be expressed as:
$\mathbf{\Sigma}$ $=\sigma_{0}^{2}\mathbf{R}+\left(\sum_{b=1}^{p}{\gamma_{b}}\mathbf{X_{b}X_{b}^{T}}\right)\odot\mathbf{V}+\sum_{b=1}^{p}{\gamma_{b}}q_{b}\mathbf{X_{b}X_{b}^{T}}$
$=\sigma_{0}^{2}\mathbf{R}+\left(\mathbf{X}\operatorname{diag}(\boldsymbol{\gamma})\mathbf{X^{T}}\right)\odot\mathbf{V}+\mathbf{X}\operatorname{diag}(\boldsymbol{\gamma}\odot\mathbf{q})\mathbf{X^{T}}$ (3)
where $\odot$ denotes elementwise multiplication, and $\operatorname{diag}(\cdot)$ denotes a diagonal matrix with the entries of the vector inside the parentheses on its diagonal. The trait values at tips can be written as:
$\mathbf{Y}=\beta_{0}\mathbf{1}+\sum_{b}\beta_{b}\mathbf{X_{b}}+\mathbf{\epsilon}$
Where $\epsilon$ follows a normal distribution with mean 0 and covariance matrix $\Sigma$, given by Equation 2.2. The main task is to select the branches that have $\beta_{b}\neq 0$ and $\gamma_{b}\neq 0$; and estimate the values of $\beta_{b}$ and $\gamma_{b}$.
3 Methods
In this section, we propose a new method to simultaneously detect the shifts in both variance and optimal values based on minimizing the loss function with L1 penalty. When $Y$ follows a multivariate normal distribution with mean and variance given by Equations 1 and 2.2, the log likelihood is
$l(\beta,\gamma)$ $=-\frac{1}{2}(\mathbf{Y}-\beta_{0}-\mathbf{X}\mathbf{\beta})^{T}\mathbf{\Sigma}^{-1}(\mathbf{Y}-\beta_{0}-\mathbf{X}\mathbf{\beta})-\frac{1}{2}\log{\det{(\mathbf{\Sigma})}}$
$=-\frac{1}{2}(\mathbf{Y}-\beta_{0}-\mathbf{X}\mathbf{\beta})^{T}\left(\sigma_{0}^{2}\mathbf{R}+\left(\mathbf{X}\operatorname{diag}(\boldsymbol{\gamma})\mathbf{X^{T}}\right)\odot\mathbf{V}+\mathbf{X}\operatorname{diag}(\boldsymbol{\gamma}\odot\mathbf{q})\mathbf{X^{T}}\right)^{-1}(\mathbf{Y}-\beta_{0}-\mathbf{X}\mathbf{\beta})$
$\qquad-\frac{1}{2}\log{\det{\left(\sigma_{0}^{2}\mathbf{R}+\left(\mathbf{X}\operatorname{diag}(\boldsymbol{\gamma})\mathbf{X^{T}}\right)\odot\mathbf{V}+\mathbf{X}\operatorname{diag}(\boldsymbol{\gamma}\odot\mathbf{q})\mathbf{X^{T}}\right)}}$
To select the shifts in optimal values ($\beta_{i}\neq 0$) and the shifts in variance ($\gamma_{i}\neq 0$), we use an L1 penalty in the loss function to conduct the feature selection as in LASSO. Therefore, the loss function to be optimized is given by:
$L(\beta_{1},...,\beta_{p},\gamma_{1},...,\gamma_{p})$ $=l(\beta_{1},...,\beta_{p},\gamma_{1},...,\gamma_{p};\textbf{Y})+\text{Penalty}$
$=\frac{1}{2}(\mathbf{Y}-\beta_{0}-\mathbf{X}\mathbf{\beta})^{T}\left(\sigma_{0}^{2}\mathbf{R}+\left(\sum_{i=1}^{p}\gamma_{i}\mathbf{X_{i}X_{i}^{T}}\right)\odot\mathbf{V}+\sum_{i=1}^{p}\gamma_{i}q_{i}\mathbf{X_{i}X_{i}^{T}}\right)^{-1}$
$\qquad(\mathbf{Y}-\beta_{0}-\mathbf{X}\mathbf{\beta})+\frac{1}{2}\log{\det{\left(\sigma_{0}^{2}\mathbf{R}+\left(\sum_{i=1}^{p}\gamma_{i}\mathbf{X_{i}X_{i}^{T}}\right)\odot\mathbf{V}+\sum_{i=1}^{p}\gamma_{i}q_{i}\mathbf{X_{i}X_{i}^{T}}\right)}}$
$\qquad+\lambda_{1}\|\boldsymbol{\beta}\|_{1}+\lambda_{2}\|\boldsymbol{\gamma}\|_{1}$ (4)
3.1 Optimization
In this paper, we do not estimate the parameter $\alpha$ jointly with other parameters, but instead treat it as fixed using an ad-hoc estimate obtained from the null model fitted by the phylolm R package. Therefore, in this section, we treat it as fixed during the model fitting process. In the simulations, we demonstrate that the estimation of $\alpha$ does not significantly affect the detection results. The primary objective here is to find the parameters $\boldsymbol{\beta}$, $\boldsymbol{\gamma}$, and $\sigma_{0}^{2}$ that minimize the loss function described in Equation 3. When $\boldsymbol{\gamma}$ is fixed, the problem reduces to a standard LASSO formulation, which can be efficiently solved.
For optimizing $\boldsymbol{\gamma}$, we employ coordinate-wise proximal gradient descent, as described by Parikh and Boyd, (2014). The proximal gradient algorithm is a powerful tool for handling non-differentiable optimization problems. It decomposes the objective function into two components: a differentiable part and a non-differentiable part, allowing for iterative updates that combine gradient descent with the proximal operator. The proximal operator serves to handle the non-differentiable component, promoting sparsity and effectively reducing the influence of irregularities in the model. By applying the proximal gradient algorithm in our optimization process, we can efficiently minimize the objective function even when the problem involves non-smooth terms, such as LASSO regularization.
The loss function for $\gamma_{k}$ can be written as:
$L(\gamma_{k})=g(\gamma_{k})+\lambda_{2}\|\gamma_{k}\|_{1}.$
The derivative of $g(\gamma_{k})$ is given by:
$\nabla g(\gamma_{k})$ $=-\frac{1}{2}\mathbf{r}^{T}\mathbf{\Sigma^{-1}}\left(\mathbf{\left(X_{k}X_{k}^{T}\right)}\odot\mathbf{V}\right)\mathbf{\Sigma^{-1}}\mathbf{r}-\frac{1}{2}q_{k}\|\mathbf{X_{k}^{T}\Sigma^{-1}r}\|^{2}+\frac{1}{2}\operatorname{tr}\left(\left(\mathbf{\left(X_{k}X_{k}^{T}\right)}\odot\mathbf{V}\right)\mathbf{\Sigma}^{-1}\right)$
$\qquad+\frac{1}{2}q_{k}\mathbf{X_{k}^{T}\Sigma^{-1}X_{k}},$
where $\mathbf{r}=\mathbf{Y}-\beta_{0}-\mathbf{X^{T}}\mathbf{\beta}$. The proximal operator is given by:
$\text{prox}_{\lambda_{2}h}(z)=S_{\lambda_{2}}(z)=\begin{cases}z-\lambda_{2}&z>\lambda_{2},\\ 0&-\lambda_{2}\leq z\leq\lambda_{2},\\ z+\lambda_{2}&z<-\lambda_{2}.\end{cases}$
The proximal algorithm here steps in the direction $\nabla g(\gamma_{k})$ but setting any values that are close to zero equal to zero. For $\sigma^{2}_{0}$, there is no penalty on the parameter. Therefore, we use gradient descent to update $\sigma_{0}^{2}$. To avoid negative values of $\sigma_{0}^{2}$, we use $\tau_{0}=\log(\sigma_{0}^{2})$ for optimization. The gradient of $\tau_{0}$ is given by:
$\frac{\partial L}{\partial\tau}$ $=\left(-\frac{1}{2}\mathbf{r^{T}}\mathbf{\Sigma}^{-1}\mathbf{R}\mathbf{\Sigma}^{-1}\mathbf{r}+\frac{1}{2}\text{tr}\left(\mathbf{R}\mathbf{\Sigma}^{-1}\right)\right)e^{\tau_{0}}.$
We initialize $\boldsymbol{\beta}$ and $\boldsymbol{\gamma}$ as $\mathbf{0}$. In each iteration, we first update the covariance matrix $\mathbf{\Sigma}$ and transform $\mathbf{X}$ and $\mathbf{Y}$ with the current $\mathbf{\Sigma}^{-1/2}$. Then we update $\boldsymbol{\beta}$ by applying LASSO on the transformed data. Then we update $\boldsymbol{\gamma}$ using the proximal algorithm in an elementwise manner. After that, we update $\tau_{0}^{2}$ using gradient descent. The algorithm is summarized in Algorithm 1.
Algorithm 1 Optimization with $\boldsymbol{\beta}$, $\boldsymbol{\gamma}$ and $\tau_{0}$
1:Inputs:
2:      $M$: maximum number of steps; $t$: step size; $\epsilon$: error tolerance; $\mathbf{V}$; $\mathbf{R}$; $\mathbf{Y}$; $\mathbf{X}$; $\lambda_{2}$
3:Initialize:
4:      $\beta_{i}^{0}\leftarrow 0$, $i=1,\ldots,p$ $\gamma_{i}^{0}\leftarrow 0$, $i=1,\ldots,p$ $\tau_{0}\leftarrow 0$ $L\leftarrow Inf$
5:for s = 1 to M do
6:     $\mathbf{\Sigma}\leftarrow e^{(\tau_{0})}\mathbf{R}+\left(\sum_{i=1}^{p}\gamma_{i}\mathbf{X_{i}X_{i}^{T}}\right)\odot\mathbf{V}+\sum_{i=1}^{p}\gamma_{i}q_{i}\mathbf{X_{i}X_{i}^{T}}$
7:     $\mathbf{Y^{\prime}}=\mathbf{\Sigma^{-1/2}Y}$; $\mathbf{X^{\prime}}=\mathbf{\Sigma^{-1/2}X}$
8:     $\beta\leftarrow LASSO(\mathbf{Y^{\prime}},\mathbf{X^{\prime}})$
9:     for k = 1 to p do
10:         calculate $\nabla g(\gamma_{k})$
11:         update $\gamma_{k}\leftarrow S_{\lambda_{2}t}(\gamma_{k}-t\nabla g(\gamma_{k}))$
12:         update $\mathbf{\Sigma}$
13:     end for
14:     calculate the gradient for $\tau_{0}$: $\frac{\partial L}{\partial\tau_{0}}$
15:     update $\tau_{0}\leftarrow\tau_{0}-t\frac{\partial L}{\partial\tau_{0}}$
16:     update the loss function $L$ with Equation 3
17:     if update for the loss function $L$ in the iteration $<\epsilon$ then
18:         break
19:     end if
20:end for
3.2 Model selection
For the model selection process, we use a strategy to fine-tune the parameters $\lambda_{1}$ and $\lambda_{2}$ to strike an optimal balance between model simplicity and performance.
  • Fixing $\lambda_{2}$: We begin by setting $\lambda_{2}$ to a fixed value, which controls the degree of regularization for variance shifts. With $\lambda_{2}$ fixed, we focus on adjusting $\lambda_{1}$, which regulates the penalization of shifts in the optimal values.
  • Cross-validation for $\lambda_{1}$: For each fixed value of $\lambda_{2}$, we use cross-validation to find the best $\lambda_{1}$. We first de-correlate the data with the estimated covariance matrix. We then use the cv.glmnet function from the R package glmnet to efficiently perform cross-validation here. We use cross-validation because cv.glmnet allows for quick and easy computation of cross-validation results, making it more convenient and efficient for selecting the optimal parameters.
  • Iterating over different $\lambda_{2}$ values: After optimizing $\lambda_{1}$ for one fixed $\lambda_{2}$, we repeat this process for multiple values of $\lambda_{2}$, producing a set of candidate models, each associated with different pairs of $\lambda_{1}$ and $\lambda_{2}$.
  • Selecting the best model using BIC: To identify the optimal model, we compare all candidate models across the range of $\lambda_{2}$ values by using the Bayesian Information Criterion (BIC). The model with the lowest BIC is selected as the final model, as it achieves the best trade-off between goodness of fit and model complexity, minimizing the risk of overfitting.
We choose BIC over other criteria based on empirical findings in Zhang et al., (2024), where pBIC was found to be a more conservative criterion. While a conservative selection method can be beneficial in some cases, detecting shifts in variance is inherently more challenging than detecting shifts in optimal values. A less conservative criterion like BIC is preferable here, as it increases sensitivity to shifts in variance while still penalizing overly complex models.
4 Simulations
4.1 Comparison of performance
We generated simulation data under an Ornstein–Uhlenbeck (OU) model with a fixed root at a value $y_{0}=\theta_{0}=0$, using the Anolis lizard phylogeny from Mahler et al., 2013b . This is an ultrametric tree and we rescaled the tree height to 1. The selection strength parameter was set to $\alpha=1$, which corresponds to a phylogenetic half-life of approximately $t_{1/2}=\ln(2)/\alpha\approx 0.693$. The stationary diffusion variance to $\sigma_{0}^{2}=2$. Each parameter setting was repeated across 50 simulation replicates. To assess the robustness of our conclusions, we also performed additional simulations with different values of $\alpha$, which are presented in the supplementary materials and yield results consistent with the main findings. In the following simulations, $\alpha$ is assumed to be unknown. For ShiVa, we estimate $\alpha$ by fitting a null model (without shifts) using the phylolm package.
We considered three simulation scenarios: (1) a single shift in the optimal value, (2) a single shift in the variance, and (3) one shift in the optimal value combined with one shift in the variance.
For the optimal value shift, we varied the shift size across the values $-5$, $-3$, $-1$, $1$, $3$, and $5$. For the variance shift scenarios, we varied the shift magnitude $\gamma_{b}$ across the values $-1.5$, $-1$, $1$, $1.5$, $2$, $3$, $5$, and $7$. In the combined scenario, we first fixed the variance shift at $\gamma_{b}=5$ and varied the optimal value shift as in Scenario (1), and then fixed the optimal value shift at $5$ and varied the variance shift as in Scenario (2). For the setting of $\gamma_{b}$, the largest magnitude considered was 7, which corresponds to a variance that is 3.5 times greater than the original value, given that the baseline $\sigma^{2}$ is 2. In the case study section, we observe examples where the estimated variance shift reaches up to 7 times the original value. Therefore, we believe that the simulated shift magnitudes fall within a realistic and meaningful range.
We compared the performance of different methods using four key metrics: True Positives (TP), False Positives (FP), predictive log-likelihood, and computational time. We count the number of correctly detected true shifts as True Positives (TP), and the number of detected shifts that do not correspond to any true shift as False Positives (FP). The positions of shifts are not always identifiable: for example, under a 3-branch tree, any two shifts can lead to an identical model. In such cases, TP and FP may not be reliable metrics, as they would count the true model as including one false positive and missing one true positive, even though the models are effectively equivalent. However for the simulation studies presented here, the true shifts are identifiable. Figure 3 illustrates the TP and FP results across the various methods under different scenarios. When there is only a shift in the optimal value, ShiVa achieves a well-balanced trade-off between True Positives and False Positives. It effectively detects shifts in the optimal value without generating excessive False Positives. The performance of ShiVa is comparable to other methods that only consider shifts in the optimal value even if only a shift in optimal value is present. In contrast, PCMFit produces a much higher number of False Positives, particularly when the signal size is small.
In the case of shifts in variance, ShiVa performs well in detecting variance shifts effectively while maintaining control over False Positives. Although PCMFit shows a higher True Positive rate compared to ShiVa, it comes at the cost of generating significantly more False Positives. Methods like $\ell$1ou and phyloEM, which are designed to detect shifts in optimal value, face challenges when a shift in variance is present. As the magnitude of the variance shift increases, these methods incorrectly interpret the variance signal as multiple shifts in the optimal value. Consequently, their False Positives increase substantially. In contrast, ShiVa successfully distinguishes between variance and mean shifts, effectively controlling the False Positive rate for optimal value shifts, even as the variance shift signal strengthens. We have more simulations with varying $\alpha$ values in the Supplementary Material and the conclusion remains similar.
Figure 4 shows the detection frequencies of false positive shifts in the optimal value when a true variance shift is present on branch 195 with shift magnitude $\gamma_{b}=7$. The results indicate that the true variance shift on branch 195 is often misinterpreted as shifts in the optimal value on its descendant branches by methods that only consider optimal value shifts, such as $\ell$1ou and phyloEM. PCMFit tends to produce false positive detections more broadly across the tree. In contrast, ShiVa, which models shifts in diffusion variance, is less likely to produce false positive detections of optimal value shifts.
When both shifts in optimal value and variance are present, ShiVa continues to perform competitively. Although it detects slightly fewer True Positives compared to PCMFit, its False Positive rate is much lower, indicating a better balance between sensitivity and specificity. PCMFit, while capable of identifying more shifts overall, tends to produce excessive False Positives. This leads to reduced reliability in shift detection, especially in noisy data.
(a) 1 shift in optimal value
(b) 1 shift in variance
(c) 1 shift in optimal value + 1 shift in variance (fixed gamma)
(d) 1 shift in optimal value + 1 shift in variance (fixed beta)
Figure 3: Comparison of True positives and False positives of different methods. Note: In panel (a, left), when $\beta=-5$ or $\beta=5$, all methods correctly detect the shift (True Positives = 1), resulting in overlapping points.
Figure 4: Detection frequencies of false positive shifts in the optimal value under a true variance shift. The pink box highlights the true variance shift on branch 195 (with shift magnitude $\gamma_{b}=7$). Branch 195 refers to the edge label number of the shifted branch.
In addition to TP and FP, we also evaluated predictive log-likelihood to assess the prediction accuracy of the models estimated by different methods. Regarding identifiability, the OU model is identifiable when no shifts are present. However, previous studies have shown that the exact number and locations of shifts along an edge cannot always be fully determined due to identifiability issues. For example, any two of the three branches connected to a node can yield equivalent models. Even if the selected shifts do not match the true model exactly, choosing a close surrogate shift may be preferable to missing the shift entirely. In such cases, the true positive versus false positive framework might misrepresent performance: a method selecting a surrogate shift is penalized with a false positive, even though it provides a more reasonable approximation than failing to detect the shift. To assess how well each method generalizes beyond the training data, we use predictive log-likelihood. Specifically, for each of the 50 training datasets, we estimate model parameters (including the locations and magnitudes of shifts) and use them to compute the expected mean vector and covariance matrix under the fitted model. We then generate 1,000 new test datasets independently using the same phylogeny and true evolutionary process as the training data. For each of these test datasets, we calculate the log-likelihood under the estimated model from the corresponding training set, and take the average to obtain a predictive log-likelihood for that training replicate. This process results in 50 predictive log-likelihood values, from which we report summary statistics (e.g., mean and median). This metric assesses how well the inferred model generalizes to new data drawn from the same evolutionary process, providing insight into the biological relevance of the estimated shifts. Since the log-likelihoods of the true models vary, we compute the predictive log-likelihood difference as the predictive log-likelihood of each estimated model minus the log-likelihood of the corresponding true model. This normalization allows for fair comparison across methods, where a higher score indicates better performance. Figure 5 illustrates the differences in predictive log-likelihood between the estimated models and the true model. When shifts occur only in optimal value, ShiVa performs comparably to $\ell$1ou and phyloEM, demonstrating similar predictive precision. However, when shifts in variance are introduced, ShiVa shows a notable improvement, achieving higher predictive log-likelihood than the other methods. In contrast, PCMFit consistently exhibits the lowest predictive log-likelihood across all scenarios, likely due to overfitting.
(a) 1 shift in optimal value
(b) 1 shift in variance
(c) 1 shift in optimal value + 1 shift in variance (fixed gamma)
(d) 1 shift in optimal value + 1 shift in variance (fixed beta)
Figure 5: Comparison of predictive log-likelihood of different methods
Finally, we compare the average computational time across different methods. Methods that only account for shifts in optimal values, such as $\ell$1ou, are the fastest, with $\ell$1ou being particularly efficient due to its use of the LASSO path method. ShiVa performs similarly to phyloEM but is slightly slower in comparison. On the other hand, PCMFit requires significantly more time due to its handling of complex computations, especially when both shifts in optimal values and variance are present. The current PCMFit implementation allows parallel computing to partially compensate for the disadvantage of the method in terms of computational complexity. For the comparisons in Table 1, we ran all methods single-threaded to provide a fairer comparison of computation required by all methods.
Table 1: Computational Time Comparison for Different Methods
Average Computation Time (seconds) Shift Position (mean) Beta Shift Position (variance) Gamma ShiVa l1ou-pBIC l1ou-BIC phyloEM PCMFit 71 1 0 0 95.38 8.09 8.31 68.15 11484.33 71 3 0 0 77.39 8.29 8.57 70.10 13592.56 71 5 0 0 73.86 8.59 8.78 69.92 12227.06 71 -1 0 0 143.39 7.86 8.24 67.81 12890.20 71 -3 0 0 75.70 8.31 8.45 70.02 16215.21 71 -5 0 0 70.33 8.37 8.57 69.35 11962.69 0 0 195 1 70.22 8.08 8.30 68.63 15171.28 0 0 195 1.5 83.31 8.08 8.42 67.72 15680.57 0 0 195 2 77.95 8.04 8.32 70.84 20622.02 0 0 195 3 87.73 8.52 8.81 70.93 16349.04 0 0 195 5 90.05 8.10 8.34 67.70 17358.66 0 0 195 7 97.68 8.18 8.53 67.85 19644.00 0 0 195 -1 101.92 8.04 8.30 67.95 15269.57 0 0 195 -1.5 102.83 7.58 7.74 65.42 15557.74 71 5 195 5 79.32 8.30 8.68 68.18 22836.10 71 3 195 5 129.45 8.55 8.73 68.68 24499.03 71 1 195 5 75.43 8.26 8.92 69.12 18023.39 71 -1 195 5 85.47 7.93 8.24 69.71 19622.12 71 -3 195 5 68.37 7.88 8.51 69.10 24580.25 71 -5 195 5 77.10 7.89 8.21 70.45 23177.43 71 5 195 1 76.73 8.23 8.53 70.06 16574.76 71 5 195 1.5 72.93 8.27 8.59 69.33 13153.55 71 5 195 2 76.58 8.18 8.61 68.71 15742.43 71 5 195 3 90.45 8.17 8.54 69.52 22800.45 71 5 195 7 87.67 8.42 8.67 69.44 23418.19 71 5 195 -1 92.83 8.23 8.45 70.31 16642.62 71 5 195 -1.5 95.15 8.31 8.51 70.79 17517.06
4.2 Relationship between shift position and detection difficulty
To investigate the difficulty of detecting shifts at specific positions on a phylogenetic tree, we conducted simulations for every internal branch of the tree. For each branch, we performed 50 simulations where a shift in the optimal value occurred and 50 simulations where a shift in variance occurred. We then calculated the detection probabilities for each shift using ShiVa, considering different positions on the tree. Figure 6 presents the detection probabilities for various branches, highlighting the different detection patterns for shifts in optima versus shifts in variance.
The results indicate that shifts in optima and shifts in variance exhibit distinct detection patterns. Shifts in the optimal value tend to be more detectable when they occur near the root or at the tips, while those occurring in the middle regions of the tree are generally harder to identify. In contrast, shifts in variance are more challenging to detect overall. Our findings show that shifts in variance are more likely to be detected when they occur on branches with a greater number of descendant tips and longer branch lengths, which amplify the signal of the variance change. Notably, the two branches immediately below the root node consistently show lower detection probabilities for variance shifts. This is not due to a lack of signal, but because shifts on these two branches are statistically non-identifiable. This observation is consistent with our analysis in the Evolutionary shifts in optimal value and variance subsection. Overall, these results highlight how the structure of the phylogeny and the position of shifts critically influence the detectability of evolutionary changes, with variance shifts requiring more informative configurations for reliable inference compared to shifts in optima.
Figure 6: The detection probability of a shift on different branches (ShiVa)
4.3 The influence of estimation of $\alpha$
We conducted simulations to examine how misestimating $\alpha$ affects ShiVa’s ability to detect shifts. The true value of $\alpha$ was set to 1 (i.e., $\log(\alpha)=0$). To assess the impact of varying estimates, we ran ShiVa using a range of fixed $\alpha$ values corresponding to $\log(\alpha)=-5,-3,-1,0,1,3,5$. We evaluated the results by comparing predictive log-likelihoods across these settings. Importantly, the predictive log-likelihood was computed using the true $\alpha$ value to ensure fair comparison of the detected shifts and to avoid confounding effects from the misestimated $\alpha$.
For this analysis, we focused on scenarios with one shift in the optimal value and one shift in variance, representing a typical use case for ShiVa. The results suggest that, in most cases, moderate misestimation of $\alpha$ has little effect on ShiVa’s performance. However, the impact becomes more pronounced in extreme cases — particularly when the estimated $\alpha$ is very large (e.g., $\exp(5)$) and the shift signal size is also very large. This indicates that while ShiVa is generally robust to $\alpha$ misestimation, severe overestimation combined with strong signals may reduce detection accuracy. Overall, this analysis supports the practical simplification of using a reasonable ad-hoc estimate of $\alpha$ in applications.
(a) 1 shift in optimal value: $\beta$ = 3
(b) 1 shift in variance: $\gamma$ = 3
Figure 7: Comparison of predictive log-likelihood of ShiVa with different estimation of $\alpha$
5 Case study
To illustrate the application of our method, we analyzed two empirical datasets. The first dataset concerns floral diameter in the parasitic plant family Rafflesiaceae, using the phylogeny flowerTree from Davis et al., (2007). The second dataset involves buccal morphology in Centrarchidae sunfishes, based on a time-calibrated phylogeny of 28 species from Revell and Collar, (2009). We applied ShiVa to log-transformed floral diameter in the Euphorbiaceae dataset and log-transformed buccal length in the Centrarchidae dataset, comparing results with $\ell$1ou+pBIC, $\ell$1ou+BIC, PhylogeneticEM, and PCMFit. Notably, ShiVa incorporates cross-validation during parameter tuning, introducing some randomness into the results. To ensure robustness, we recommend running ShiVa three times and selecting the model with the lowest BIC, as we did in our empirical analyses.
Figure 8 presents the shift detection results on the flower dataset. ShiVa identified one shift in the optimal value and one shift in evolutionary variance, with estimated values of $\sigma^{2}=0.33$ and $\gamma=2.42$, respectively. Compared to the results of $\ell$1ou+pBIC, ShiVa detects one additional variance shift. The shift size of 2.42 is substantial, especially when considered relative to the original variance, highlighting a significant change in evolutionary dynamics.
Although ShiVa appears to yield a lower log-likelihood and higher BIC than $\ell$1ou+pBIC, this comparison includes the estimation of the selection strength parameter $\alpha$. When comparing detection results under a fixed $\alpha$, ShiVa achieves a higher log-likelihood and lower BIC than $\ell$1ou+pBIC, suggesting that modeling variance shifts improves model fit. Notably, compared to $\ell$1ou+pBIC, ShiVa detects an additional variance shift on branch 16. To evaluate the significance of this shift, we performed a likelihood ratio test. The log-likelihood of the model with only a mean shift on branch 41 is $-21.8940$, while the model that also includes a variance shift on branch 16 achieves a log-likelihood of $-18.3639$. The resulting test statistic corresponds to a $p$-value of $0.0293$, which is below the $0.05$ threshold, indicating that the variance shift on branch 16 is statistically significant. PhyloEM and $\ell$1ou+BIC yield identical shift detections and attain the highest log-likelihood among all methods.
To better understand the behavior of each method, we grouped the detected models into four representative types: (1) ShiVa: one shift in optimal value and one shift in variance; (2) $\ell$1ou+pBIC: one shift in optimal value; (3) $\ell$1ou+BIC / PhyloEM: three shifts in optimal value; (4) PCMFit: changes in all parameters on two shifted branches. To assess how well each method performs when a given model is the true generative process, we simulate 100 datasets from each estimated model and re-apply all five methods to the simulated data. We then compute the predictive log-likelihood to compare model fit across methods. As shown in Table 2, ShiVa consistently achieves relatively high and stable predictive log-likelihood, even when the data are generated from models estimated by other methods. This suggests that ShiVa produces robust and reasonable detection results across a variety of underlying evolutionary scenarios, supporting its credibility and generalizability.
Figure 8: The detection results on the flower dataset using different methods are shown. Green circles indicate shifts in optimal trait values, while pink squares denote shifts in evolutionary variance. The numbers inside the markers correspond to the node indices where shifts occur, and the colored numbers next to them represent the estimated magnitudes of those shifts.
Table 2: The median predictive log-likelihood of different methods, evaluated on datasets simulated from each of the estimated models. (flower dataset)
Model Simulated ShiVa $\ell$1ou+pBIC $\ell$1ou+BIC phyloEM PCMFit True
(1) -23.8826 -28.8382 -34.8087 -28.6572 -167.6036 -18.3726
(2) -21.2532 -23.0154 -29.6700 -23.0154 -153.1015 -18.0685
(3) -20.0300 -21.1384 -18.1560 -20.6989 -234.6207 -7.0669
(4) -23.2938 -26.3329 -70.2521 -26.1872 -284.8035 -13.4989
Figure 9 summarizes the results from the In the Centrarchidae analysis, ShiVa detected two shifts in the optimal trait value and three shifts in evolutionary variance. It estimated a saturated model by assigning an extremely small value to the original $\sigma^{2}$. Among the three variance shifts, the one on branch 48—with a magnitude of 0.309—is the most substantial, while the other two primarily restore typical levels of diffusion variance. $\ell$1ou+pBIC identified two shifts in the optimal trait value, while $\ell$1ou+BIC inferred 11 such shifts. PhylogeneticEM and PCMFit detected no shifts under their respective model selection criteria.
ShiVa achieved a higher log-likelihood than $\ell$1ou+pBIC, PhylogeneticEM, and PCMFit. Although $\ell$1ou+BIC attained the highest log-likelihood, it did so by detecting a large number of shifts in the optimal value—including several within the clade descending from the branch where ShiVa detected a variance shift. For example, ShiVa detected a variance shift on branch 48 with a substantial magnitude of 0.309. Under that branch, $\ell$1ou+BIC detected two shifts in the optimal value (branches 29 and 31), and $\ell$1ou+pBIC also detected a shift at branch 29. Despite its higher log-likelihood, $\ell$1ou+BIC resulted in a higher BIC than ShiVa.
Using this dataset, we grouped the estimated models into four representative types: (1) ShiVa: 2 shifts in optimal value and 3 shifts in variance; (2) $\ell$1ou+pBIC: 2 shifts in optimal value; (3) $\ell$1ou+BIC: 11 shifts in optimal value; (4) PhyloEM / PCMFit: no shifts detected.
As in the previous analysis, we simulated 100 datasets from each estimated model and re-applied all five methods to the simulated data. We then computed the predictive log-likelihood to evaluate model fit across methods. As shown in Table 3, ShiVa consistently achieves relatively high and stable predictive log-likelihoods, except for model (2). Its performance is affected in scenarios with a large number of shifts in the optimal value.
Figure 9: The detection results on the centrarchidae data using different methods are shown. Green circles indicate shifts in optimal values, and pink squares indicate shifts in evolutionary variance. The numbers inside the markers represent the branch indices where shifts occur, while the colored numbers next to them denote the estimated shift magnitudes.
Table 3: The median predictive log-likelihood of each method, evaluated on simulated datasets generated from each estimated model. (sunfish dataset)
Model Simulated ShiVa $\ell$1ou+pBIC $\ell$1ou+BIC phyloEM PCMFit True
(1) 27.7391 17.7871 2.2102 19.3929 19.9311 72.3057
(2) 43.2906 49.1707 43.7983 41.2296 40.9689 57.3213
(3) 25.9643 49.8254 85.6678 39.0104 46.2593 99.8206
(4) 41.8720 41.3245 31.9973 43.1736 44.7510 46.0488
6 Conclusion
In this article, we have used a multi-optima multi-variance OU process to describe an evolutionary process where abrupt shifts can occur in either optimal value or variance. We then proposed a new method to simultaneously detect shifts in optimal value and shifts in variance. We implemented the method in R. Our package is available from the first author’s GitHub page https://github.com/WenshaZ/ShiVa. Furthermore, we have conducted simulation studies to show the effectiveness of our method to detect both kinds of shifts and compared it to methods which only detect shifts in optimal value.
Our results showed that ShiVa effectively balances True Positives (TP) and False Positives (FP) across different scenarios. When there was only a shift in the optimal value, ShiVa performed comparably to other methods, such as $\ell$1ou and phyloEM, maintaining similar predictive accuracy while achieving a better balance between TP and FP compared to PCMFit, which tended to overfit in weak signal scenarios. When a shift in variance was present, ShiVa demonstrated a significant advantage over other methods, achieving higher predictive log-likelihood and effectively distinguishing between variance and mean shifts. This led to fewer False Positives compared to $\ell$1ou and phyloEM, which often misinterpreted variance shifts as multiple shifts in optimal value. Although PCMFit achieved higher TP, it generated significantly more FP, indicating overfitting.
Our empirical analyses on floral diameter in Euphorbiaceae and buccal morphology in Centrarchidae sunfishes demonstrate the practical utility of ShiVa. In both datasets, ShiVa successfully identified meaningful variance shifts that were missed or misinterpreted by methods focusing only on optimal value shifts. While ShiVa may yield slightly lower likelihoods in some cases, it achieves better model fit when accounting for variance shifts and maintains strong predictive performance across models. These results underscore ShiVa’s robustness and its ability to capture complex evolutionary dynamics that may be overlooked by existing approaches.
For model simplicity, we assume that the variance parameter is constant along each branch and that all shifts in variance happen at the beginning of the branch. Simulations of cases where these assumptions are violated show that our method is robust to misspecification in these assumptions. An interesting future research direction is to extend our model to allow shifts in variance at internal positions along a branch. This would allow us to estimate the exact time of a shift in variance. Writing the likelihood for this case is straightforward, but more work may be needed to ensure stability of the estimates. Another direction for future work is to improve the computational efficiency of the method. The likelihood calculation involves a large number of matrix inverse computations,which can be computationally expensive for large trees. Bastide et al., (2021) provide an efficient algorithm to calculate the log likelihood and its derivatives for certain phylogenetic models. If this method could be adapted to our method, it would lead to a major improvement in the computational speed, allowing our method to scale to larger phylogenies.
7 Data accessibility
The phylogenetic tree of Anolis lizards is provided by Mahler et al., 2013a and it can be accessed via the R package l1ou (https://github.com/khabbazian/l1ou). The phylogenetic tree and trait data for the Euphorbiaceae flower dataset (Davis et al.,, 2007) are available through the R package phylolm, and the Centrarchidae sunfish dataset (Revell and Collar,, 2009) can be accessed via the R package phytools. Both packages are available on CRAN. Our R package ShiVa is available at https://github.com/WenshaZ/ShiVa, and the simulation and case study code can be found at https://github.com/WenshaZ/ShiVa-Experiments.
8 Author contributions
All authors contributed to conceive the ideas, design the methodology and the simulations. WZ conducted the simulations, analyzed the results and wrote the manuscript under the guidance of LSTH and TK. LSTH and TK reviewed and edited the manuscript.
9 Acknowledgement
LSTH was supported by the Canada Research Chairs program, the NSERC Discovery Grant RGPIN-2018-05447, and the NSERC Discovery Launch Supplement. TK was supported by the NSERC Discovery Grant RGPIN-2023-03332.
Supplemental Materials
A The effect of shift position within a branch
In Trait evolution with shifts in both optimal value and variance, we noted that our assumptions about the shifts in variance—at most one shift will occur on any branch; and the shifts occur at the beginning of the branches—are restrictions on the space of possible models, so in particular if these assumptions are not satisfied, then the model is misspecified. In this subsection, we conduct simulations in which these assumptions are violated.
Firstly, we conduct a series of simulations with the shift occuring in different positions along a fixed branch. The different locations of the shift are shown in Figure 10 (left). Figure 11 shows the True Positives v.s. False Positives and the difference of predictive log-likelihood between ShiVa and the true model with shifts in different locations. Location here is a parameter ranging from 0 to 1, with 0 meaning the beginning of the branch and 1 meaning the end of the branch. The ability of the method to detect the shift in variance is not greatly impacted by the position of the shift along the branch. Furthermore, when the shift size is small, the log-likelihood of the model is not much reduced in situations where the location of the shift is not at the beginning of the branch. However, when the magnitude of the shift is larger, the model misspecification does cause a substantial decrease in log-likelihood when the shift is not located at the begining of the branch. Overall, our method is robust to violations of our assumption that shifts occur at the beginning of a branch.
Figure 10: Left: The different locations of shifts on the same branch; Right: Two opposite shifts occur on the same branch
Figure 11: The true positive v.s. false positive and the predictive log-likelihood difference of ShiVa with shifts at different locations
Secondly, we simulate scenarios with two shifts on a single branch. If shifts are both positive or both negative, the data will show a stronger signal than when just one shift occurs. Therefore, we simulate situations where two opposite shifts occur (at locations = 0 and 1) (Figure 10 right). To compare, we also simulate situations where only one shift occurs at the beginning of that branch. Table 4 shows that when the two opposite shifts occur on that branch, our method cannot accurately detect the shift in variance on that branch and the false positive number of shifts in optimal value increases. In this case, the violation of the assumption causes some difficulties for our method. However, in the general case, it is not a major concern.
Table 4: The performance of ShiVa when two opposite shifts occur on the same branch
Shift Size 20 15 10 5 1
False positive(optimal value) opposite shifts 0.66 0.63 0.49 0.48 0.44
single shift 0.4 0.5 0.26 0.49 0.43
True positive(variance) opposite shifts 0 0 0 0 0
single shift 0.9 0.86 0.84 0.41 0.01
False positive(variance) opposite shifts 0.01 0 0.02 0 0
single shift 0.73 0.7 0.57 0.21 0.02
B Supplementary figures and tables
(a) 3 shifts in optimal value
(b) 7 shifts in optimal value
(c) 1 shift in variance (different position)
(d) 2 shifts in variance
Figure 12: Comparison of True positives and False positives of different methods (Supplementary simulations)
(a) 3 shifts in optimal value
(b) 7 shifts in optimal value
(c) 1 shift in variance (different position)
(d) 2 shifts in variance
Figure 13: Comparison of predictive log-likelihood (difference with true model) of different methods (Supplementary simulations)
Table 5: Computational Time Comparison for Different Methods (Supplementary simulations)
Average Computation Time (seconds) Shift Position (mean) Beta Shift Position (variance) Gamma ShiVa l1ou-pBIC l1ou-BIC phyloEM PCMFit 43,71,192 1 0 0 86.80 7.66 8.43 69.64 11561.58 43,71,192 3 0 0 74.95 8.37 8.36 68.09 27477.04 43,71,192 5 0 0 88.87 8.76 8.86 68.05 36900.89 43,71,192 -1 0 0 71.88 7.99 8.40 69.21 13952.25 43,71,192 -3 0 0 67.89 8.61 8.56 70.58 35867.14 43,71,192 -5 0 0 76.54 9.13 9.16 72.69 40202.81 6,71,197,3,88,191,98 1 0 0 117.94 7.80 8.05 70.53 14863.06 6,71,197,3,88,191,98 3 0 0 80.94 6.06 6.08 70.63 31135.83 6,71,197,3,88,191,98 5 0 0 81.68 8.55 8.86 67.94 59322.66 6,71,197,3,88,191,98 -1 0 0 83.20 7.79 8.04 69.02 15652.61 6,71,197,3,88,191,98 -3 0 0 98.53 7.42 6.84 69.47 29763.69 6,71,197,3,88,191,98 -5 0 0 69.55 7.99 7.85 70.15 60076.49 0 0 196 1 68.06 8.08 8.35 69.46 14515.10 0 0 196 1.5 72.20 8.03 8.31 69.51 15649.65 0 0 196 2 83.65 8.39 8.54 70.18 17266.36 0 0 196 3 90.27 7.61 8.14 69.24 18907.37 0 0 196 5 89.08 8.27 8.42 69.61 18364.71 0 0 196 7 101.44 8.03 8.51 67.09 18858.03 0 0 196 -1 99.09 8.26 8.66 67.89 16773.86 0 0 196 -1.5 113.54 8.21 8.72 67.30 19419.61 0 0 195,130 1,-0.5 80.75 8.37 8.68 70.94 16275.18 0 0 195,130 3,-1 93.39 8.11 8.39 70.21 18297.14 0 0 195,130 5,-1.5 159.13 12.06 12.58 116.55 36602.57 0 0 195,130 7,-1.5 114.15 8.60 8.86 69.08 22939.48 0 0 195,130 1 85.99 8.15 8.31 70.17 15881.60 0 0 195,130 3 91.66 8.10 8.38 70.03 17969.21 0 0 195,130 5 104.25 8.69 9.17 71.10 25327.97 0 0 195,130 7 95.08 8.63 8.94 69.17 23545.64
(a) $\alpha$ = 0.01
(b) $\alpha$ = 0.1
(c) $\alpha$ = 10
Figure 14: Comparison of True positives and False positives of different methods (Varying $\alpha$)